# Multi-Player, Multi-Strategy Quantum Game Model for Interaction-Aware Decision-Making in Autonomous Driving

### v1

@@ II-C Quantum Game Theory
For a more detailed comparison between CGT and QGT, the reader can refer to [46, 37]. After the brief overview of the conceptual differences between CGT and GQT, we discuss relevant studies that have applied quantum game theory.
@@ III-D Quantum Models
  • Quantum Gates $QG_{i}$, which corresponds to specific values of $\theta_{i}$. These gates provide discrete strategic choices, offering simpler control of the qubits compared to a unitary operator. We defined $QG_{i}=\{H,\sigma_{x},\sigma_{y},\sigma_{z},I_{2}\}$, where:
@@ IV Results
TABLE I: Quantum parameters used for QGDM-U and QGDM-Q models. 2P / 2S denotes games with two players and two strategies. EPD stands for Equal Probability Distribution. The initial quantum state $\big|\psi_{0}\big|^{2}=[0,0,1,0]^{T}$ indicates that the game starts in the joint action profile where player 1 selects $a_{1}^{1}$ and player 2 selects $a_{2}^{0}$, this corresponds, for example, the ego decelerating and the interacting vehicle accelerates in the merging use case with two players.
@@
TABLE II: Quantitative results obtained for the highway scenario. Compared with baseline methods. $\mathbf{n_{col}}$: number of collisions detected, HD: mean Headway Distance (between EV and vehicle ahead), $\mathbf{\bar{v}_{EV}}$: mean ego speed, $\mathbf{\bar{a}_{EV}}$: mean ego acceleration, $\mathbf{\bar{t}}$: mean duration of an episode, $\mathbf{\rho_{CLL}}$: proportion of left lane changes maneuvers, $\mathbf{\rho_{CLR}}$: proportion of right lane changes maneuvers, and $\mathbf{\rho_{KL}}$: proportion of keep lanes maneuvers.
@@ IV-C Highway
We also evaluated our approach in highway scenarios. Here, the EV can choose among $\mathcal{A}_{EV}=\{\text{Change Lane Left},\text{Change Lane Right},\text{Idle}\}$, while the IV can select $\mathcal{A}_{IV}=\{\text{Accelerate},\text{Decelerate},\text{Idle}\}$. The IV acting as the opponent player in the game is selected as the vehicle closest to the EV. The longitudinal behavior of the ego is controlled by the IDM model [43], with parameters set to represent a regular driver profile [30]. To account for more than two strategies per player, we employ the quantum circuit depicted in subfigure (c) of Figure 4(a). For each episode, the initial lane of the ego vehicle is randomly selected among the 4 available ones. The results from this simulation are shown in Table II. Here, we noticed several notable differences. First of all, the average time $\mathbf{\bar{t}}$ to complete an episode is lowest for the IDMxMOBIL method, which is not surprising as this model is specifically designed to optimize efficiency on highways, and other road users are also modeled using IDMxMOBIL. COR-MP, the utility-based method, is the most conservative, with a mean headway distance HD of 85.46 meters. Overall, the quantum models did not produce any collisions in this scenario, demonstrating their applicability even in regular highway conditions where interactions are less intense compared to roundabout or merging situations. QGDM-U exhibits a higher $\mathbf{\bar{t}}$ compared to QGDM-G and IDMxMOBIL due to performing additional lane changes ($\rho_{CLL}$ and $\rho_{CLR}$), which occasionally causes the ego vehicle to become temporarily blocked behind another vehicle and unable to change lanes anymore.

### v2

# Multi-Player, Multi-Strategy Quantum Game Model for Interaction-Aware Decision-Making in Automated Driving

@@ II-C Quantum Game Theory
For a more detailed comparison between CGT and QGT, the reader can refer to [46, 37]. After the brief overview of the conceptual differences between CGT and QGT, we discuss relevant studies that have applied quantum game theory.
@@ III-D Quantum Models
  • Quantum Gates $QG_{i}$, which correspond to specific values of $\theta_{i}$. These gates provide discrete strategic choices, offering simpler control of the qubits compared to a unitary operator. We defined $QG_{i}=\{H,\sigma_{x},\sigma_{y},\sigma_{z},I_{2}\}$, where:
@@ IV Results
TABLE I: Quantum parameters used for QGDM-U and QGDM-Q models. 2P / 2S denotes games with two players and two strategies. EPD stands for Equal Probability Distribution. The initial quantum state $\big|\psi_{0}\big|^{2}=[0,0,1,0]^{T}$ indicates that the game starts in the joint action profile where player 1 selects $a_{1}^{1}$ and player 2 selects $a_{2}^{0}$, corresponding, for example, to the ego decelerating and the interacting vehicle accelerating in the merging use case with two players.
@@
TABLE II: Quantitative results obtained for the highway scenario, compared against baseline methods. $\mathbf{n_{col}}$: number of collisions detected, HD: mean Headway Distance (between EV and vehicle ahead), $\mathbf{\bar{v}_{EV}}$: mean ego speed, $\mathbf{\bar{a}_{EV}}$: mean ego acceleration, $\mathbf{\bar{t}}$: mean duration of an episode, $\mathbf{\rho_{CLL}}$: proportion of left lane changes maneuvers, $\mathbf{\rho_{CLR}}$: proportion of right lane changes maneuvers, and $\mathbf{\rho_{KL}}$: proportion of keep lanes maneuvers.
@@ IV-C Highway
We also evaluated our approach in highway scenarios. Here, the EV can choose among $\mathcal{A}_{EV}=\{\text{Change Lane Left},\text{Change Lane Right},\text{Idle}\}$, while the IV can select $\mathcal{A}_{IV}=\{\text{Accelerate},\text{Decelerate},\text{Idle}\}$. The IV acting as the opponent player in the game is selected as the vehicle closest to the EV. The longitudinal behavior of the ego is controlled by the IDM model [43], with parameters set to represent a regular driver profile [30]. To account for more than two strategies per player, we employ the quantum circuit depicted in subfigure (c) of Figure 4(a). For each episode, the initial lane of the ego vehicle is randomly selected among the 4 available ones. The results from this simulation are shown in Table II. Here, we noticed several notable differences. First of all, the average time $\mathbf{\bar{t}}$ to complete an episode is lowest for the IDMxMOBIL method, which is not surprising as this model is specifically designed to optimize efficiency on highways, and other road users are also modeled using IDMxMOBIL. COR-MP, the utility-based method, is the most conservative, with a mean headway distance HD of 85.46 meters. Overall, the quantum models did not produce any collisions in this scenario, demonstrating their applicability even in regular highway conditions where interactions are less intense compared to roundabout or merging situations. QGDM-U exhibits a higher $\mathbf{\bar{t}}$ compared to QGDM-G and IDMxMOBIL due to performing additional lane changes ($\rho_{CLL}$ and $\rho_{CLR}$), which occasionally cause the ego vehicle to become temporarily blocked behind another vehicle and unable to change lanes anymore.